OpenStudy (anonymous):

HELP!! fx=ln(1-lnx) need f' then solve for critical values..

OpenStudy (anonymous):

\[-\frac{ 1 }{ 1-\ln(x) }\frac{ 1 }{ x }=\frac{ df(x) }{ dx }\]

OpenStudy (anonymous):

Great, thanks..How would you solve this for critical values?

OpenStudy (anonymous):

\[-\frac{ 1 }{ x-xln(x) }=\frac{ df(x) }{ dx }\] Wheatever values of x make the function undefined (x-xln(x)=0) will be your critical values for x

OpenStudy (anonymous):

so it looks like the only value is x=0

OpenStudy (anonymous):

Because "0" is in the denominator?

OpenStudy (anonymous):

no because when x=0 the denominator is undefined. you cannot divide a number by 0

OpenStudy (anonymous):

or yes I guess that is what you were asking so yes

OpenStudy (anonymous):

Thanks for the help!!

OpenStudy (anonymous):

you're welcome